OpenStudy (anonymous):

lim as x --> 1 ((sq rt of x) -1)/(x-1)

OpenStudy (anonymous):

use conjugates

OpenStudy (anonymous):

\[ \lim_{x\rightarrow 1} \frac{\sqrt{x}-1}{x-1} = \lim_{x\rightarrow 1} \frac{(\sqrt{x}-1)(\sqrt{x}+1)}{(x-1)(\sqrt{x}+1)} \]

OpenStudy (anonymous):

\[ = \lim_{x\rightarrow 1} \frac{x-1}{(x-1)(\sqrt{x}+1)} = \lim_{x\rightarrow 1} \frac{1}{\sqrt{x}+1} = \frac{1}{2} \]

OpenStudy (anonymous):

the answer is -1

OpenStudy (anonymous):

Nope, thats wrong.

OpenStudy (anonymous):

The answer is 1/2, check if you copied the problem correctly.

OpenStudy (jamesj):

The answer is definitely 1/2. I prefer the way YTheManifold has done it, because it is very clear. But if we use a theorem and calculate it using L'Hopital's Rule, we would arrive at \[\lim_{x \rightarrow 1} \ (1/(2\sqrt{x})) / 1 = 1/2\]